What is the EAR of a 6-month account that pays an APR of 4.3%?
Answer #1

Effective annual rate is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.0430/2)^2 -1

        = 1.0435 -1

        = 0.0435*100

        = 4.35%.

Therefore, the effective annual rate of a 6 month account is 4.35%.